The sum of the numbers A and B is 116. A is 3 less than C but 4 more than B. Find the value of C?
Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39631)

Transcribing what you said into symbolism,
A+B=116; A=-3+C; A=4+B.

You have three equations. Two of the equations are statements giving expressions for A. You can use A=4+B to substitute into the sum_116 equation and solve for B.
Try to work through the rest of a solution process on your own.

Answer by lwsshak3(11628)

The sum of the numbers A and B is 116. A is 3 less than C but 4 more than B. Find the value of C?
***
A+B=116
..
A=C-3
C=A+3
..
A=B+4
B=A-4
..
A+B=116
A+A-4=116
2A=120
A=60
..
C=A+3
C=63